ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: not given Todorov, Kaltscho, Bulgaria, Aslansky, M. On the n-generated Quaternion Groups and the Hadamard Matrices The problem of Hadamard matrices is formulated as well the Williamson’s method has been given for the construction of the Hadamard matrices of higher order. Based on the author’s research over n-generated quaternion group is exposed other possibility of Hadamard matrices of higher order construction.